A calorimeter is used to try To determine the enthalpy of combustion of a candle. The calorimeter contains 7.50 x 10^2 mL of water. The initial mass of the candle is 42.3g and the initial temperature of the water is 18.1 C. After being allowed to burn for a period of time, the candle is re-measured to have a mass of 38.8g and the temperature of the water is 55.6C.
a) how much heat energy was transferred into the water ?
b) how much energy was transferred per gram of paraffin wax?
c) How much energy was transferred per mile of paraffin wax?
d) If the theoretical enthalpy of combustion of paraffin wax is -42kj/g, what is the percentage difference ?
1
Expert's answer
2019-07-15T07:04:59-0400

a) To calculate heat trransferred to water we should use the formula:

q = cmΔT

where c (for water) = 4.184 J/mol*K

m - mass of water

m = d*V = 1g/ml *750 ml = 750 g

ΔT = 55.6 - 18.1 = 37.5 0C

so, q = 4.184*750*37.5 =117675 J = 118 kJ

b) Calculate the mass of candle which was burned:

m = 42.3 - 38.3 = 4

Energy transferrred per garm of paraffin wax is :

ΔH =-q/m = -118 kJ/4 g = -29.5 kJ/g

c) Calculate moles of wax paraffin (C25H52)

n = m/M = 4/352.68 = 0.0113 mol

ΔH = -q/n = -118 kJ/0.0113 mol = -10442 kJ/mol

d) Percentage difference = |ΔH(theoretical)| - |ΔH(actual)|/ |ΔH(theoretical)| *100% = |-42|-|-29.5|/|-42| *100% = 29.8%.